OpenStudy (anonymous):

find the real numbers A and B if the line 8x-y=3 is tangent to the graph of f(X)= Ax^2+Bx+1 at x=1

OpenStudy (freckles):

Well have you found f' yet?

OpenStudy (freckles):

find f' and find the slope of (8x-y=3) set them equal you have one equation in terms of A and B. You know the line (8x-y=3) and f(x)=Ax^2+Bx+1 share at least one common point which is the point of tangency (1,f(1)) You can find f(1) by solving 8(1)-y=3 for y.

OpenStudy (anonymous):

okay thanks i will try that

OpenStudy (freckles):

@Mathboy23 how are you doing?

OpenStudy (freckles):

\[f'(1)=slope_{of}(8x-y=3) \\ f(1)=A(1)^2+B(1)+1=solve_{for y when x=1}(8x-y=3)\] You will have a system of equations to solve in the end

OpenStudy (anonymous):

I ended up getting B=0 and A=4, do you know if thats right

OpenStudy (freckles):

ok so what did you get the slope of the line 8x-y=3 was 8?

OpenStudy (anonymous):

yes i got 8 for the slope

OpenStudy (freckles):

\[f'(1)=8 \\\] and when you solve 8x-y=3 for y when x=1 you got y=5? so you have \[f(1)=5\]

OpenStudy (anonymous):

yes

OpenStudy (freckles):

so good so far

OpenStudy (freckles):

what system did you end up solving?

OpenStudy (freckles):

Like what the system of equations look like

OpenStudy (anonymous):

A+B+1=5 and 2A+B=8 then i multiplied the first equation by 2 and subtracted the second equation

OpenStudy (freckles):

Awesome your solution looks great

OpenStudy (freckles):

Do you have any questions?

OpenStudy (anonymous):

no, thank you so much for all your help :)

OpenStudy (freckles):

np